I got 4 for this equation. Is it correct.
-4x-10x=-56

Answer by stanbon(75887)

I got 4 for this equation. Is it correct.
-4x-10x=-56
-----
Substitute 4 in place of "x" and see if you get a true statement.
--------------------------------
-4*4 -10*4 = -56
-16 - 40 = -56
-56 = -56
-----------------------
True, so 4 is a solution
================================
Cheers,
Stan H.